OpenStudy (anonymous):

substitution method. 3x+3y=33

OpenStudy (ageta):

Let's solve for x. 3x+3y=33 Step 1: Add -3y to both sides. 3x+3y+−3y=33+−3y 3x=−3y+33 Step 2: Divide both sides by 3.

OpenStudy (ageta):

x=−y+11

TheSmartOne (thesmartone):

We nee two formulas.